A stone is thrown vertically upward with a velocity of 14m/s.

(a) What is the maximum height will it attain?

(b) What is its velocity on striking the ground?


1
Expert's answer
2021-03-02T07:44:39-0500

(a) Let's first find the tme that the stone takes to reach the maximum height:


"v_y=v_{0y}-gt,""0=v_{0y}-gt,""t=\dfrac{v_{0y}}{g}=\dfrac{14\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=1.43\ s."

Finally, we can find the maximum height attained by the ball:


"y_{max}=v_{0y}t-\dfrac{1}{2}gt^2,""y_{max}=14\ \dfrac{m}{s}\cdot1.43\ s-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(1.43\ s)^2=10\ m."

b) We can find the velocity of the stone as it strikes the ground from the kinematic equation:


"v_f^2=v_i^2+2gh_{max},""v_f=\sqrt{v_i^2+2gh_{max}},""v_f=\sqrt{0+2\cdot9.8\ \dfrac{m}{s^2}\cdot10\ m}=14\ \dfrac{m}{s}."